# A Forward Reachability Perspective on Control Barrier Functions and Discount Factors in Reachability Analysis

### v3

@@ 7.2 Proof of Proposition 2
$V_{\gamma}(x_{1})\leq\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}J_{\gamma}(x_{1},\mathrm{u}^{-},\hat{\xi}_{d}^{-})+\varepsilon,\vspace{-0.5em}$
@@
$V_{\gamma}(x_{1})\leq J_{\gamma}(x_{1},\mathrm{u}^{-},\hat{\xi}_{d}^{-})+\varepsilon\vspace{-0.5em}$
@@
$V_{\gamma}(x_{2})\geq\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}J_{\gamma}(x_{2},\mathrm{u}^{-},\hat{\xi}_{d}^{-})\geq J_{\gamma}(x_{2},\hat{\mathrm{u}}^{-},\hat{\xi}_{d}^{-})-\varepsilon.\vspace{-0.5em}$
@@
$=\sup_{t\in(-\infty,0]}\!\!e^{\gamma t}h_{S}(\mathrm{x}_{1}^{-}(t))\;-\!\!\sup_{t\in(-\infty,0]}\!\!e^{\gamma t}h_{S}(\mathrm{x}_{2}^{-}(t))+2\varepsilon,\vspace{-0.5em}$
@@
$\sup_{t\in(-\infty,0]}e^{\gamma t}h_{S}(\mathrm{x}_{1}^{-}(t))\leq e^{\gamma\hat{t}}h_{S}(\mathrm{x}_{1}^{-}(\hat{t}))+\varepsilon,\vspace{-0.5em}$
@@
$\leq L_{h_{S}}e^{\gamma\hat{t}}e^{-L_{f}\hat{t}}\lVert x_{1}-x_{2}\rVert+3\varepsilon\leq L_{h_{S}}\lVert x_{1}-x_{2}\rVert+3\varepsilon,\vspace{-0.5em}$
@@ 7.3 Proof of Theorem 2
$V_{\gamma}(x)=\sup_{\xi_{d}\in\Xi_{d}}\inf_{\mathrm{u}\in\mathcal{U}}\sup_{t\in[0,\infty)}e^{-\gamma t}h_{S}(\mathrm{x}(t)),$ (37)
@@
(Dynamic Programming principle [38, Lemma 3]). Suppose $\gamma>0$. For $x\in\mathbb{R}^{n}$,
[…]
@@
Next, Theorem 2 holds by [38, Lemma 3], since $V_{\gamma}$ rewritten as (37) is the unique viscosity solution to
[…]
@@ 7.4 Proof of Theorem 3
$l^{i}(\xi_{d},\mathrm{u},x)\coloneqq e^{-\gamma T}V^{i}(\mathrm{x}(-T)),\vspace{-0.5em}$
@@
$B_{T}[V^{i}]=\sup_{\xi_{d}\in\Xi_{d}}\inf_{\mathrm{u}\in\mathcal{U}}\max\{l(\xi_{d},\mathrm{u}),l^{i}(\xi_{d},\mathrm{u})\}.\vspace{-0.5em}$
[…]
@@
$\;\leq 2\varepsilon+e^{-\gamma T}\max_{x\in\mathbb{R}^{n}}\lvert V^{1}(x)-V^{2}(x)\rvert\vspace{-1em}$
@@
$\lVert B_{T}[V^{1}]-B_{T}[V^{2}]\rVert_{L^{\infty}(\mathbb{R}^{n})}\leq e^{-\gamma T}\lVert V^{1}-V^{2}\rVert_{L^{\infty}(\mathbb{R}^{n})}\vspace{-0.5em}$
Since $V_{\gamma}$ is a fixed-point solution for all $T>0$, the Banach’s contraction mapping theorem [15, Chapter 9.2] implies that $V_{\gamma}$ is the unique fixed-point solution to $B_{T}[V_{\gamma}](x)=V_{\gamma}(x)$ for all $T>0$. In addition, we have
$\lVert B_{T}[V]-V_{\gamma}\rVert_{L^{\infty}(\mathbb{R}^{n})}\leq e^{-\gamma T}\lVert V-V_{\gamma}\rVert_{L^{\infty}(\mathbb{R}^{n})}\vspace{-0.5em}$
@@ 7.5 Computation methods for $V_{\gamma}$
First, the following lemma presents a finite-horizon HJ equation for the computation of $V_{\gamma}$ using the Bellman operator $B_{T}$.
@@
$\inf_{\xi_{d}\in\xi_{d}}\sup_{\mathrm{u}}\min_{s\in[t,0]}e^{-\gamma(s-t)}(-h_{S}(\mathrm{x}(s))),\vspace{-0.5em}$
@@
[…]
$e^{-\gamma\delta}W^{+}(t+\delta,\mathrm{x}(t+\delta))\bigg{\}}.\vspace{-0.5em}$
@@
$W^{+}(0,x)=-\max\{h_{S}(x),V_{0}(x)\}\quad\text{on }\{t=0\}\times\mathbb{R}^{n},$
@@
$\quad\quad\quad\frac{\partial W^{+}}{\partial t}+\max_{u}\min_{d}\frac{\partial W^{+}}{\partial x}\cdot(-f)(x,u,d)-\gamma W^{+}(t,x)\Big{\}},\vspace{-1em}$
@@
Theorem 3 also enables other numerical schemes that are based on time discretization, like value iteration, to produce an accurate solution of $V_{\gamma}$. The following corollary of Theorem 3 provides the guarantee that the value iteration with any initial guess of $V^{0}\in\textnormal{BUC}(\mathbb{R}^{n})$ will converge to $V_{\gamma}$ with a Q-linear convergence rate specified by (44). For a given time step size $\Delta t$, the semi-Lagrangian approximation can be applied to the exact Bellman backup operator in (22) for its numerical approximation, and the resulting value function will converge to $V_{\gamma}$ when $\Delta t\rightarrow 0$ [2].
@@ 7.6 Proof of Proposition 4
(i) At $x\in\mathrm{Int}(S)$ where $V_{\gamma}$ is differentiable, the FRT-HJ-VI (21) implies that $K_{\gamma}$ is non-empty.
@@
$0\!=\!\min\Big{\{}\!\ V^{\pi}_{\gamma}(x)-h_{S}(x),\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma V^{\pi}_{\gamma}(x)\Big{\}}.$
For each $x\in\mathbb{R}^{n}$, $\min\{y-h_{S}(x),\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma y\}$ is monotonically increasing in $y$, so the equation (45) has a unique solution. Also, from the FRT-HJ-VI (21),
The last inequality holds since $V_{\gamma}-h_{S}\geq 0$ from (21) and $\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma V_{\gamma}(x)\geq 0$ since $\pi(x)\in K_{\gamma}(x)$. Equation (46) and the uniqueness of (45) imply $V_{\gamma}\equiv V^{\pi}_{\gamma}$ for any $\pi$. By replacing $V_{\gamma}$ by $V^{\pi}_{\gamma}$ in (45),
$0\!=\!\min\Big{\{}\!\ V^{\pi}_{\gamma}(x)\!-\!h_{S}(x),\min_{d}\frac{\partial V^{\pi}_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma V^{\pi}_{\gamma}(x)\Big{\}}$.
@@ 7.7 Proof of Theorem 4
    $0\leq\min\Big{\{}v(x_{0})-h(x_{0}),\max_{u\in U}\min_{d\in D}\frac{\partial v}{\partial x}(x_{0})\cdot f(x_{0},u,d)+\gamma v(x_{0})\Big{\}}.$
@@
    $0\geq\min\Big{\{}v(x_{0})-h(x_{0}),\max_{u\in U}\min_{d\in D}\frac{\partial v}{\partial x}(x_{0})\cdot f(x_{0},u,d)+\gamma v(x_{0})\Big{\}}.$
Case 1. $V_{\gamma}(x_{0})\!=\!h(x_{0})\!>\!0$: By the continuity of $h$, there exists $\varepsilon>0$ such that $V_{\gamma}(y)\!=\!h(y)$ for all $y\!\in\!B_{\varepsilon}(x_{0})$. Thus, the gradient of $V_{\gamma}$ at $x_{0}$ exists: $\frac{\partial V_{\gamma}}{\partial x}(x_{0})=\frac{\partial h}{\partial x}(x_{0})$, so for any $v$ such that $V_{\gamma}-v$ has either a local minimum or a local maximum at $x_{0}$, $\frac{\partial v}{\partial x}(x_{0})=\frac{\partial h}{\partial x}(x_{0})$. From (26),
$\max_{u\in U}\min_{d\in D}\frac{\partial v}{\partial x}(x_{0})\cdot f(x_{0},u,d)+\gamma v(x_{0})=0.\vspace{-0.5em}$ (47)
@@
Case 2. $V_{\gamma}(x_{0})\!=\!0\!>\!h(x_{0})$: By the continuity of $h$, there exists $\varepsilon>0$ such that $V_{\gamma}(y)=0$ for all $y\in B_{\varepsilon}(x_{0})$. This implies that the gradient of $V_{\gamma}$ at $x_{0}$ is $0\in\mathbb{R}^{n}$, so for any $v$ such that $V_{\gamma}-v$ has either a local minimum or a local maximum at $x_{0}$, $\frac{\partial v}{\partial x}(x_{0})=0$. Thus, (47) holds. Therefore, Statements 1) and 2) hold in this case.
Case 3. $V_{\gamma}(x_{0})\!=\!0\!=\!h(x_{0})$: From $v(x_{0})-h(x_{0})\!=\!0$, it is trivial that 2) holds, and we focus on the proof of 1). Since $V_{\gamma}-v$ has a local minimum at $x_{0}$, $\frac{\partial v}{\partial x}(x_{0})\in\partial^{-}V_{\gamma}(x_{0})$, where $\partial^{-}V_{\gamma}(x_{0})$ is the sub-differential, which is determined as $\partial^{-}V_{\gamma}(x_{0})=\text{conv}\left(\{0\}\cup\Big{\{}\frac{\partial h}{\partial x}(x_{0})\Big{\}}\right)$, where conv is a convex-hull operator. Thus, $\frac{\partial v}{\partial x}(x_{0})=\alpha\frac{\partial h}{\partial x}(x_{0})$ for some $\alpha\in[0,1]$. Thus, from (26) and $v(x_{0})=0$, (47) holds and therefore, 1) holds.
7.8 1D example for comparison of methods in Table 1
We consider a simple one-dimensional system:
$\dot{\mathrm{x}}(t)=\mathrm{x}(t)+\mathrm{u}(t),\;\mathrm{x}(0)=x,\vspace{-0.5em}$ (48)
with $u\in U=[-1,1]$, and the state domain $x\in[0,\infty)$. We do not introduce disturbance for simplicity. We consider the set $S=[0,2]$, and choose $h_{S}$ as $h_{S}(x)=\max\{2\!-\!x,-2\}$. Basically, it is a distance function cut off at the absolute value of $2$ (Figure 5 grey). We compare the results of the three backward reachability formulations studied in the previous literature [17, 2, 38, 13, 35] with our forward reachability formulation, as summarized in Table 1. We use $\gamma=2$ in all formulations. Note that the chosen $S$ is not control invariant, thus, the example is chosen not for the reachability analysis of control invariant sets, but to study the boundedness, continuity, and the solution uniqueness of the resulting value functions. The maximal control invariant set contained in $S$, is $[0,1]$, since at $x=1$, $\dot{x}$ can be maintained 0 by selecting the saturated control input $u=-1$, but for every $x$ exceeding 1, $\dot{x}>0$ for any admissible $u\in U$.
Figure 5: Various reachability value functions, summarized in Table 1, for the 1D example in (48). Only the formulations with the discounted factor (magenta and blue), including the one proposed in this work produce value functions that are bounded and continuous.
Backward Reachable Tube (BRT) without discount factor [17, 20]:
$V(x):=\underset{{\xi_{d}\in\Xi_{d}}}{\inf}\;\underset{\mathrm{u}\in\mathcal{U}}{\sup}\;\underset{t\in[0,\infty)}{\inf}h_{S}(\mathrm{x}(t)),\vspace{-0.5em}$ (49)
which characterizes the viability kernel of $S$ as $\{x|V(x)\geq 0\}$, which can be seen in Figure 5 (purple). The value function is discontinuous at $x=1$. Moreover, the corresponding HJ-PDE, given as
$0=\min\Big{\{}\ h_{S}(x)-V(x),\max_{u}\min_{d}\frac{\partial V}{\partial x}\cdot f(x,u,d)\Big{\}}$
admits non-unique solutions, for instance, $V(x)\equiv-2$ in this example is also a valid viscosity solution to (50).
BRT with discount factor [2, 38]:
$V(x):=\underset{{\xi_{d}\in\Xi_{d}}}{\inf}\;\underset{\mathrm{u}\in\mathcal{U}}{\sup}\;\underset{t\in[0,\infty)}{\inf}e^{-\gamma t}h_{S}(\mathrm{x}(t)).\vspace{-0.5em}$ (51)
In this case, as can be seen in Figure 5 (green), the value function is continuous and bounded. However, the value function is flat inside the viability kernel, which is characterized as $\{x|V(x)=0\}$.
Infinite horizon CBVF [13, 35]:
$V(x):=\underset{{\xi_{d}\in\Xi_{d}}}{\inf}\;\underset{\mathrm{u}\in\mathcal{U}}{\sup}\;\underset{t\in[0,\infty)}{\inf}e^{\gamma t}h_{S}(\mathrm{x}(t)).\vspace{-0.5em}$ (52)
Notice the flip of the sign in the factor of the exponential term, compared to (51). This formulation results in an HJ-PDE whose differential inequality matches the form of the barrier constraint. However, it results in discontinuity and unboundedness of the value function, as can be seen in Figure 5 (orange).
Our formulation: The value function is defined in (18), which is bounded and continuous, as can be seen in Figure 5 (blue). Also, this formulation admits a unique solution to the corresponding HJ-PDE in (21), and the differential inequality in the PDE matches the form of the barrier constraint. Note that in this example, $\mathrm{FRT}(\mathrm{Int}(S))$ is $[0,\infty)$, thus, in Figure 5, $V(x)>0$ everywhere.

### v4

@@ 8.5 Proof of Proposition 4
$V_{\gamma}(x_{1})\leq\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}J_{\gamma}(x_{1},\mathrm{u}^{-},\hat{\xi}_{d}^{-})+\varepsilon,\vskip-5.0pt$
@@
$V_{\gamma}(x_{1})\leq J_{\gamma}(x_{1},\mathrm{u}^{-},\hat{\xi}_{d}^{-})+\varepsilon\vskip-5.0pt$
@@
$V_{\gamma}(x_{2})\geq\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}J_{\gamma}(x_{2},\mathrm{u}^{-},\hat{\xi}_{d}^{-})\geq J_{\gamma}(x_{2},\hat{\mathrm{u}}^{-},\hat{\xi}_{d}^{-})-\varepsilon.\vskip-5.0pt$
@@
$=\sup_{t\in(-\infty,0]}\!\!e^{\gamma t}h_{S}(\mathrm{x}_{1}^{-}(t))\;-\!\!\sup_{t\in(-\infty,0]}\!\!e^{\gamma t}h_{S}(\mathrm{x}_{2}^{-}(t))+2\varepsilon,\vskip-5.0pt$
@@
$\sup_{t\in(-\infty,0]}e^{\gamma t}h_{S}(\mathrm{x}_{1}^{-}(t))\leq e^{\gamma\hat{t}}h_{S}(\mathrm{x}_{1}^{-}(\hat{t}))+\varepsilon,\vskip-5.0pt$
@@
$\leq L_{h_{S}}e^{\gamma\hat{t}}e^{-L_{f}\hat{t}}\lVert x_{1}-x_{2}\rVert+3\varepsilon\leq L_{h_{S}}\lVert x_{1}-x_{2}\rVert+3\varepsilon,\vskip-5.0pt$
@@ 8.6 Proof of Theorem 3
$V_{\gamma}(x)=\sup_{\xi_{d}\in\Xi_{d}}\inf_{\mathrm{u}\in\mathcal{U}}\sup_{t\in[0,\infty)}e^{-\gamma t}h_{C}(\mathrm{x}(t)),$ (50)
@@
(Dynamic Programming principle [49, Lemma 3]). Suppose $\gamma>0$. For $x\in\mathbb{R}^{n}$,
$V_{\gamma}(x)\!=\!\!\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}\!\!\max\!\Big\{$ $\max_{t\in[-T,0]}e^{\gamma t}h_{C}\left(\mathrm{x}^{-}(t)\right),$ (51)
@@
Next, Theorem 3 holds by [49, Lemma 3], since $V_{\gamma}$ rewritten as (50) is the unique viscosity solution to
$0=\min\big\{$ $V_{\gamma}(x)-h_{C}(x),$
@@ 8.7 Proof of Theorem 4
$l^{i}(\xi_{d},\mathrm{u},x)\coloneqq e^{-\gamma T}V^{i}(\mathrm{x}(-T)),\vskip-5.0pt$
@@
$B_{T}[V^{i}]=\sup_{\xi_{d}\in\Xi_{d}}\inf_{\mathrm{u}\in\mathcal{U}}\max\{l(\xi_{d},\mathrm{u}),l^{i}(\xi_{d},\mathrm{u})\}.\vskip-5.0pt$
Without loss of generality, let $B_{T}[V^{1}](x)\geq B_{T}[V^{2}](x)$. For any $\varepsilon>0$, $\exists\bar{\xi}_{d}$ such that $B_{T}[V^{1}]-\varepsilon<\inf_{\mathrm{u}}$ $\max\{l(\bar{\xi}_{d},\mathrm{u}),l^{1}(\bar{\xi}_{d},\mathrm{u})\}$, and $\exists\bar{\mathrm{u}}$ such that $\inf_{\mathrm{u}}\max\{$ $l(\bar{\xi}_{d},\mathrm{u}),l^{2}(\bar{\xi}_{d},\mathrm{u})\}\!+\!\varepsilon\!>\!\max\{l(\bar{\xi}_{d},\bar{\mathrm{u}}),l^{2}(\bar{\xi}_{d},\bar{\mathrm{u}})\}$. Then,
@@
$\;\leq 2\varepsilon+e^{-\gamma T}\max_{x\in\mathbb{R}^{n}}\lvert V^{1}(x)-V^{2}(x)\rvert\vskip-10.00002pt$
@@
$\lVert B_{T}[V^{1}]-B_{T}[V^{2}]\rVert_{L^{\infty}(\mathbb{R}^{n})}\leq e^{-\gamma T}\lVert V^{1}-V^{2}\rVert_{L^{\infty}(\mathbb{R}^{n})}\vskip-5.0pt$
Since $V_{\gamma}$ is a fixed-point solution for all $T>0$, the Banach’s contraction mapping theorem [19, Chapter 9.2] implies that $V_{\gamma}$ is the unique fixed-point solution to $B_{T}[V_{\gamma}](x)=V_{\gamma}(x)$ for all $T>0$. In addition, we have
$\lVert B_{T}[V]-V_{\gamma}\rVert_{L^{\infty}(\mathbb{R}^{n})}\leq e^{-\gamma T}\lVert V-V_{\gamma}\rVert_{L^{\infty}(\mathbb{R}^{n})}\vskip-5.0pt$
@@ 8.8 Computation methods for $V_{\gamma}$
Theorem 4 allows the computation of $V_{\gamma}$ with a finite-horizon HJ-PDE, which is guaranteed to converge to $V_{\gamma}$ as $T\rightarrow\infty$. Also, the theorem enables other numerical schemes based on time discretization, such as value iteration.
First, the following lemma presents a finite-horizon HJ equation for the computation of $V_{\gamma}$.
@@
$\inf_{\xi_{d}\in\xi_{d}}\sup_{\mathrm{u}}\min_{s\in[t,0]}e^{-\gamma(s-t)}(-h_{S}(\mathrm{x}(s))),\vskip-5.0pt$
@@
$W^{+}(t,x)\!=\!\inf_{\xi_{d}\in\Xi_{d}}\sup_{\mathrm{u}\in\mathcal{U}}\min\bigg\{$ $\!\min_{s\in[t,t+\delta]}e^{-\gamma(s-t)}(-h_{S}(\mathrm{x}(s))),$
$e^{-\gamma\delta}W^{+}(t+\delta,\mathrm{x}(t+\delta))\bigg\}.\vskip-5.0pt$
@@
$W^{+}(0,x)=-\max\{h_{S}(x),V_{0}(x)\}\qquad\text{on }\{t=0\}\times\mathbb{R}^{n},$
@@
$\qquad\qquad\qquad\frac{\partial W^{+}}{\partial t}+\max_{u}\min_{d}\frac{\partial W^{+}}{\partial x}\cdot(-f)(x,u,d)-\gamma W^{+}(t,x)\Big\},\vskip-9.24994pt$
@@
Theorem 4 also enables other numerical schemes that are based on time discretization, like value iteration, to produce an accurate solution of $V_{\gamma}$. The following corollary of Theorem 4 provides the guarantee that the value iteration with any initial guess of $V^{0}\in\textnormal{BUC}(\mathbb{R}^{n})$ will converge to $V_{\gamma}$ with a Q-linear convergence rate specified by (57). For a given time step size $\Delta t$, the semi-Lagrangian approximation can be applied to the exact Bellman backup operator in (22) for its numerical approximation:
$V^{k+1}(x)=\max\{h_{S}(x),\min_{u\in U}\max_{d\in D}e^{-\gamma\Delta t}V^{k}(x-\Delta tf(x,u,d))\},$
and the resulting value function will converge to $V_{\gamma}$ when $\Delta t\rightarrow 0$ [2].
@@ 8.9 Proof of Proposition 5
(i) At $x\in\mathrm{FRT}(\mathrm{Int}(C))$ where $V_{\gamma}$ is differentiable, the HJ-FRT-VI (21) implies that $K_{\gamma}$ is non-empty.
@@
$0\!=\!\min\Big\{\!\ V^{\pi}_{\gamma}(x)-h_{C}(x),\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma V^{\pi}_{\gamma}(x)\Big\}.$
For each $x\in\mathbb{R}^{n}$, $\min\{y-h_{C}(x),\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma y\}$ is monotonically increasing in $y$, so the equation (58) has a unique solution. Also, from the HJ-FRT-VI (21),
$0\!=\!\min\!\Big\{\!\ V_{\gamma}(x)\!-\!h_{C}(x),\max_{u}\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,u,d)+\gamma V_{\gamma}(x)\Big\}\geq$
$\min\!\Big\{\!\ V_{\gamma}(x)\!-\!h_{C}(x),\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma V_{\gamma}(x)\Big\}\!\geq\!0.$
The last inequality holds since $V_{\gamma}-h_{C}\geq 0$ from (21) and $\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma V_{\gamma}(x)\geq 0$ since $\pi(x)\in K_{\gamma}(x)$. Equation (59) and the uniqueness of (58) imply $V_{\gamma}\equiv V^{\pi}_{\gamma}$ for any $\pi$. By replacing $V_{\gamma}$ by $V^{\pi}_{\gamma}$ in (58),
$0\!=\!\min\Big\{\!\ V^{\pi}_{\gamma}(x)\!-\!h_{C}(x),\min_{d}\frac{\partial V^{\pi}_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma V^{\pi}_{\gamma}(x)\Big\}$.
@@ 8.10 Proof of Theorem 5
    $0\leq\min\Big\{v(x_{0})-h_{S}(x_{0}),\max_{u\in U}\min_{d\in D}\frac{\partial v}{\partial x}(x_{0})\cdot f(x_{0},u,d)+\gamma v(x_{0})\Big\}.$
@@
    $0\geq\min\Big\{v(x_{0})-h_{S}(x_{0}),\max_{u\in U}\min_{d\in D}\frac{\partial v}{\partial x}(x_{0})\cdot f(x_{0},u,d)+\gamma v(x_{0})\Big\}.$
Case 1. $V_{\gamma}(x_{0})\!=\!h_{S}(x_{0})\!>\!0$: By the continuity of $h_{S}$, there exists $\varepsilon>0$ such that $V_{\gamma}(y)\!=\!h_{S}(y)$ for all $y\!\in\!B_{\varepsilon}(x_{0})$. Thus, the gradient of $V_{\gamma}$ at $x_{0}$ exists: $\frac{\partial V_{\gamma}}{\partial x}(x_{0})=\frac{\partial h_{S}}{\partial x}(x_{0})$, so for any $v$ such that $V_{\gamma}-v$ has either a local minimum or a local maximum at $x_{0}$, $\frac{\partial v}{\partial x}(x_{0})=\frac{\partial h_{S}}{\partial x}(x_{0})$. From (26),
$\max_{u\in U}\min_{d\in D}\frac{\partial v}{\partial x}(x_{0})\cdot f(x_{0},u,d)+\gamma v(x_{0})=0.\vskip-5.0pt$ (60)
@@
Case 2. $V_{\gamma}(x_{0})\!=\!0\!>\!h_{S}(x_{0})$: By the continuity of $h_{S}$, there exists $\varepsilon>0$ such that $V_{\gamma}(y)=0$ for all $y\in B_{\varepsilon}(x_{0})$. This implies that the gradient of $V_{\gamma}$ at $x_{0}$ is $0\in\mathbb{R}^{n}$, so for any $v$ such that $V_{\gamma}-v$ has either a local minimum or a local maximum at $x_{0}$, $\frac{\partial v}{\partial x}(x_{0})=0$. Thus, (60) holds. Therefore, Statements 1) and 2) hold in this case.
Case 3. $V_{\gamma}(x_{0})\!=\!0\!=\!h_{S}(x_{0})$: From $v(x_{0})-h_{S}(x_{0})\!=\!0$, it is trivial that 2) holds, and we focus on the proof of 1). Since $V_{\gamma}-v$ has a local minimum at $x_{0}$, $\frac{\partial v}{\partial x}(x_{0})\in\partial^{-}V_{\gamma}(x_{0})$, where $\partial^{-}V_{\gamma}(x_{0})$ is the sub-differential, which is determined as $\partial^{-}V_{\gamma}(x_{0})=\text{conv}\left(\{0\}\cup\Big\{\frac{\partial h_{S}}{\partial x}(x_{0})\Big\}\right)$, where conv is a convex-hull operator. Thus, $\frac{\partial v}{\partial x}(x_{0})=\alpha\frac{\partial h_{S}}{\partial x}(x_{0})$ for some $\alpha\in[0,1]$. Thus, from (26) and $v(x_{0})=0$, (60) holds and therefore, 1) holds.